The Lopez-Portillo Company has $11.2 million in assets, 60 percent financed by debt and 40 percent financed by common stock. The interest rate on the debt is 9 percent and the par value of the stock is $10 per share. President Lopez-Portillo is considering two financing plans for an expansion to $21 million in assets. Under Plan A, the debt-to-total-assets ratio will be maintained, but new debt will cost a whopping 12 percent! Under Plan B, only new common stock at $10 per share will be issued. The tax rate is 40 percent. a. If EBIT is 10 percent on total assets, compute earnings per share (EPS) before the expansion and under the two alternatives. (Round your answers to 2 decimal places.) Earnings per Share Current Plan B b. What is the degree of financial leverage under each of the three plans? (Round your answers to 2 decimal places.) Degree of Financial Leverage Current Plan A Plan B c. If stock could be sold at $20 per share due to increased expectations for the firms sales and earnings, what impact would this have on earnings per share for the two expansion alternatives? Compute earnings per share for each. Round your answers to 2 decimal places.) Earnings er Share Plan A Plan B
Answer #1

Solution:-

(a)

EPS (Current)= {(EBIT-Interest)*(1-tax)}/no. of shares

EPS (Current)= [{(11.2*10%) - (11.2*60%*9%)}*(1-40%)] / {(11.2*40%)/10}

EPS (Current)= [(1.12 - 0.6048)*(0.6)] / 0.448

EPS (Current)= $0.69 per share

EPS (Plan A)= {(EBIT-Interest)*(1-tax)}/no. of shares

EPS (Plan A)= [{(21*10%) - (11.2*60%*9%) - (9.8*60%*12%)}*(1-40%)] / {(21*40%)/10}

EPS (Plan A)= [{(2.1 - 0.6048 - 0.7056)}*(0.6)] / 0.84

EPS (Plan A)= $0.564 per share

EPS (Plan B)= {(EBIT-Interest)*(1-tax)}/no. of shares

EPS (Plan B)= [{(21*10%) - (11.2*60%*9%)}*(1-40%)] / [{(11.2*40%)+9.8}/10]

EPS (Plan B)= [(2.1 - 0.6048)*(0.6)] / 1.428

EPS (Plan B)= $0.63 per share

(b)

DFL (Current) = EBIT/EBT = (11.2*10%) / {(11.2*10%) - (11.2*60%*9%)} = 1.12 / 0.5152 = 2.17 times

DFL (Plan A) = EBIT/EBT = (21*10%) / {(21*10%) - (11.2*60%*9%) - (9.8*60%*12%)} = 2.1 / 0.7896 = 2.66 times

DFL (Plan B) = EBIT/EBT = (21*10%) / {(21*10%) - 0.6048} = 2.1 / 1.4952 = 1.4 times

(c)

EPS (Plan A)= {(EBIT-Interest)*(1-tax)}/no. of shares

EPS (Plan A)= [{(21*10%) - (11.2*60%*9%) - (9.8*60%*12%)}*(1-40%)] / [{(11.2*40%)/10}+{(9.8*40%)/20}]

EPS (Plan A)= [(2.1 - 0.6048 - 0.7056)*(0.6)] / 0.644

EPS (Plan A)= $0.74 per share

EPS (Plan B)= {(EBIT-Interest)*(1-tax)}/no. of shares

EPS (Plan B)= [{(21*10%) - (11.2*60%*9%)}*(1-40%)] / [{(11.2*40%)/10}+{(9.8)/20}]

EPS (Plan B)= [(2.1 - 0.6048)*(0.6)] / 0.938

EPS (Plan B)= $0.956 per share